Instrument Log

 

NFOV     CF      Lamont, OK
 

Date

Technician

Comments

6/14/07

MK

The NFOV needs to be removed from service.Powered off and removed system at 2000 GMT.

10/19/06

MK

The 2-Channel NFOV needs to be removed to be recalibrated before deployment to AMF. Removed 2-channel NFOV at 1900 GMT.

4/13/06

PAD

Notified by computer department the 2-channel NFOV is sending the same file repeatedly to SDS system.  Requested instrument power cycled.  Power cycled instrument and data logger at 1941 GMT.  RTS at 1956 GMT.

3/23/06

MK

The 2-channel NFOV has returned and needs to be reinstalled. Installed NFOV and checked communications. Data collections were enabled.

11/16/05

MR

The NFOV single channel has arrived on site and needs to be installed. Reinstalled the *NFOV.  RTS at 1824 GMT.  *New - 93900

5/6/05

MR

The instrument mentor wants the NFOV pulled and sent back to him due to a failed heater circuit.  Shut down and removed the instrument at 1900 GMT.  Will return to the mentor.  RTS at 2015

2/9/05

MR

The instrument mentor wants the 2-channel NFOV removed and shipped back to him for upgrades.  Removed the instrument.  Returned the instrument to the mentor.  RTS at 2055 GMT.

2/8/05

PAD

Condensation observed inside window.  Removed window and cleaned condensation.  RTS at 1538 GMT.

2/7/05

PAD

Condensation observed inside window.  Removed window and cleaned condensation.  RTS at 1516 GMT.

2/2/05

PAD

Small amount of condensation observed inside window.  Removed window and cleaned condensation.  RTS at 1537 GMT.

2/1/05

PAD

Small amount of condensation observed inside window.  Removed window and cleaned condensation.  RTS at 1531 GMT.

1/31/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1558 GMT.

1/27/05

PAD

Condensation observed inside window.  Removed window and cleaned condensation.  RTS at 1534 GMT.

1/26/05

PAD

Condensation observed inside window.  Removed window and cleaned condensation.  RTS at 1523 GMT.

1/24/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1539 GMT.

1/21/05

PAD

Window has small amount of condensation inside.  Removed window and cleaned condensation.  RTS at 1631 GMT.

1/21/05

PAD

Window has small amount of condensation inside.  Removed window and cleaned condensation.  RTS at 1631 GMT.

1/20/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1546 GMT.

1/19/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1538 GMT.

1/18/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1743 GMT.

1/13/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1539 GMT.

1/12/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1554 GMT.

1/11/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1540 GMT.

1/10/05

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1619 GMT.

1/7/05

PAD

Condensation observed inside window.  Removed window and cleaned condensation.  RTS at 1550 GMT.

1/6/05

PAD

Condensation observed inside window.  Removed window and cleaned condensation.  RTS at 1642 GMT.

12/17/04

MR

The instrument mentor called and notified me that the single channel NFOV tube heater seemed to not be working.  T/S found that the tube heater was warm to the touch.  Observed the following readings on the tube heater channel; 1390 and 2445.  Reseated the heater cable.  Will report the readings to the mentor and await further instructions.  RTS at 2045.

11/24/04

PAD

Small amount of condensation found on inside of window.  Removed and cleaned window.  RTS at 1600 GMT.

11/2/04

PAD

Window has small amount of condensation inside.  Removed and cleaned window.  RTS at 1558 GMT.

10/29/04

PAD

Window has condensation inside.  Removed window and cleaned condensation.  RTS at 1446 GMT.

10/28/04

PAD

NFOV window has a very small amount of condensation inside.  Removed window and cleaned condensation from inside.  RTS at 1420 GMT.

10/26/04

PAD

NFOV window has a very small amount of condensation inside.  Removed window and cleaned condensation from inside.  RTS at 1545 GMT.

10/22/04

PAD

Window has small amount of condensation on inside.  Removed window and cleaned condensation.  RTS at 1418 GMT.

10/21/04

MR

The instrument mentor requests that the NFOV be removed and checked for water inside the sensor.  Removed the sensor at 1830 GMT.  T/S found that water had deposited a film on the sensor window inside the sensor tube.  Cleaned the window.  The rest of the sensor checked OK.  Replaced the sensor at 2005 GMT.  RTS at 2015.

10/20/04

PAD

Window has condensation inside.  Removed window and cleaned condensation.  Cannot leave window off to let instrument dry out, as humidity is high.  Installed window.  RTS at 1432 GMT.

10/19/04

PAD

Window has condensation inside.  Removed window and cleaned.  Replaced window.  RTS at 1607 GMT.

10/18/04

PAD

Window has condensation inside.  Removed window at 1410 GMT to let instrument dry out.  Installed window at 2130 GMT.  RTS  2145 GMT.

10/15/04

PAD

Window has condensation inside.  Removed window at 1452 GMT to let instrument dry out.  Installed window at 2130 GMT.  RTS 2145 GMT.

10/14/04

MR

The NFOV sensor tube is leaking during rain storms.  Removed and resealed the sensor tube glass window.  RTS at 1940

10/14/04

PAD

Window has condensation on inside.    Removed window and cleaned condensation.  RTS at 1505 GMT.

10/13/04

PAD

At 1610 GMT on 10/12/04, condensation was observed inside window.  Removed window at 1610 GMT to let instrument dry out.  Replaced window on 10/13/04 at 1420 GMT.  RTS at 1435 GMT.         

9/27/04

PAD

Window has condensation inside.  Removed window at 1430 GMT and cleaned condensation.  Replaced window and RTS at 1445 GMT.

9/24/04

PAD

NFOV 2 channel (2nd unit) has condensation inside window.  Removed and cleaned window at 1605 GMT.   RTS at 1620 GMT.

9/24/04

PAD

Window has condensation inside.  Removed and cleaned window at 1418 GMT.  Removed window so instrument could dry out at 1442 GMT.  Replaced window at 1603 GMT.  RTS at 1618 GMT.

9/23/04

PAD

Condensation has formed on inside of window.  Removed window at 1444 GMT and dried condensation.  Replaced window and RTS at 1459 GMT.

9/15/04

MR

Instrument mentor is on site to reinstall the NFOV.  The installation of the NFOV started on 9/15/05 and was completed at 1730 GMT on 9/16/04.  Notified the computer department that collections were ready to be enabled.  Communications with the instrument where checked and they checked OK.  Instrument mentor checked the data and the data checked OK.  Also installed the NFOV 2 channel next to the NFOV for inter-comparison.  RTS at 1730 GMT.

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power  was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

6/17/02

MR

The instrument housing was found to be extremely hot this morning.  T/S found that the instrument heater circuit had shorted out and caused damage to the instrument.  Notified the mentor and he recommended that we shut the system down until next week when he is expected to be onsite to perform repairs.  Shut the system down at 1815 GMT.

6/17/02

CT

The instrument had condensation inside the window. The instrument was very hot. The problem was reported to Mike Rainwater. Troubleshooting is in progress. RTS at 1529 GMT.

6/14/02

MR

Power restored.  System operational.

6/13/02

MR

The site has lost power due to storms in the area at 1415 GMT.  The instrument is shut down at this time.

6/3/02

CT

The velcro around the insulated padding needed to be replaced. There was no velcro available. The mentor was contacted and he gave his permission to use duct tape in place of the velcro. The duct tape was applied to the insulated padding at 1745 GMT. RTS at 1800 GMT.

3/19/02

SS

The SDS indicated no data files. The instrument was power cycled at 1610 GMT.  RTS 1625 GMT.  Status was OK at this time.

2/12/02

MR

The AC power to the site was lost on 1/30/02 due to local ice storm.  The instrument was found to have not been powered up on 2/11/02.  T/S found that the breaker for the AC power had tripped.  Reset the breaker.  AC power was restored.  Status ok.

11/20/01

MR

The power to the site will be shut off for about 1 hour due to a site water leak repair. Instrument was shut down from 1500 to 1600 GMT.  RTS at 1615 GMT.

8/20/01

MR

The power was lost to the site on 8/15/01 from 0030 GMT to 0400 GMT.  This instrument needs to be power cycled to reestablish communications with the SDS. Power cycled the instrument at 1425 GMT.  RTS at 1440 GMT.

12/19/00

MR

Instrument mentor is onsite to perform T/S on the NFOV tube heater circuit. T/S on the tube heater showed nothing wrong with the electronics. Cleaned and reseated the connector and the tube heater began to work better but was not working at maximum efficiency. Mentor suspects that the connections may not be able to handle the current load and may be degrading.

12/12/00

CAM

Due to a hard drive failure on the SDS Development system, data stopped being collected from the NFOV at 2152 GMT on 12/5/00. The hard drive was replaced and normal data collections resumed at 1839 GMT on 12/12/00 after all the necessary configurations were set.

5/23/00

MR

John Schmelzer sent a new resistor to replace the gain resistor in channel 1 on the data logger. Powered off the NFOV at 1908. R/R the resistor. Powered up the instrument at 1911 GMT. RTS at 1915.

2/29/2000

CAM

The NFOV was removed by the mentor on 4/23/99 for calibration. Mentor returned the NFOV and installed on 2/29/2000. SDS personnel had to make some changes to the collection and ingest processes due to the instrument being down so long prior to Y2K. Collections were successful starting at 2214 GMT on 3/1.

4/23/99

JT

The instrument mentor was on site this week and removed the NFOV sensor for calibration.  This instrument will be out of service until it returns.

4/8/99

CAM

The level bubble not entirely within the inner circle. Leveled instrument at 1352 GMT.

3/15/99

MW

The CF lost power due to a snow storm on 3/14/99 at 0030 GMT. The NFOV has been down since that time. Powered the system back up after AC power was restored. RTS at 2030 GMT. Status OK.

11/30/98

MW

The instrument mentor requested that the course temperature adjustment potentiometer be turned 1/4 turn clockwise. Adjusted the course temperature potentiometer 1/4 turn clockwise at 1735 GMT.

11/30/98

MW

The instrument mentor indicated that no files have been collected from the NFOV since the collection of 11/23/98 at 0830 GMT. All subsequent collections failed starting with the 11/24/98 0231 GMT collection. Checked the complete communications path from the SDS to the instrument. No problems found. Attempted to establish communications directly with the instrument via laptop computer. The system would not respond to operator commands. Checked the instrument power supply with a voltmeter and the meter indicated 0 Vdc. Flipped the power supply toggle switch off and checked the power supply fuse. The fuse was open. Replaced the fuse and flipped the power supply toggle switch back on. Checked the power supply voltage again with a voltmeter and the meter indicated 13.5 Vdc. Hooked the laptop computer up to the logger and was able to receive a 'hello' back. Was also able to receive a 'hello' back from the system remotely from Dev10. RTS at 1745 GMT.

11/18/98

MW

The instrument mentor asked site operations to perform a minor adjustment to the sensor head heater. Adjusted the Prop Gain Adjustment potentiometer one full turn clockwise and adjusted the Fine Temperature Adjustment potentiometer 1/4 turn clockwise at 1842 GMT as per mentor request.

11/5/98

MW/MK

The SDS had not collected a data file from the NFOV since 11/2/98 at 0244 GMT. The COMLOG indicated that the SDS could not get the attention of the datalogger. Troubleshooting efforts revealed a broken fiber connector in the junction box located on the ground next to the MWR. Moved the NFOV fibers to another set of connectors in the junction box and moved the NFOV fiber jumpers in the East end of the Optical Trailer to the appropriate connectors in the patch panel. Labeled the new location as NFOV. The SDS successfully collected a file at 0426 GMT on 11/6/98.

10/29/98

MK

The SDS has not received any data files from the NFOV. Power cycled the NFOV at 1818 GMT. RTS at 1834 GMT.

7/27/98

Mentor

Instrument installed by mentor week of 7/27/98. Daily PM checks started on 8/3/98.